| Xtb . dx Targe Itano tano da costa
Answer #1

slayton e la J xub der X +b [i +52 Who Now Had fi= 1 de Reet g-x² = +² -2x dezatot 3 xde=-tdt. then i tot 7. 4 - -t tc, F1 =Ez = b. fint() + Cz. then I = fi + f2 =-19-22 + a + b sin () +C --Jg-ek 76 fint ($) + G=C2 I =-1g-p? + b fin ) +C1 where. GI altidtil caso. fed =1.] Is 7€ +C. loq TT Is are to - Where c is constant